# The Role of Human Creativity in the Presence of AI Creativity Tools at Work: A Case Study on AI-Driven Content Transformation in Journalism

### v1

@@ 4.4.1. Learning the Limits of the Tool, and Reasserting Professional Standards
Early news reels were problematic because the news conveyed in them was unclear. This was likely due to a combination of problems: being too long, having meaningless dialog lines, and not containing enough facts. Additionally, when the tone was too positive, the reels felt like ads, not news, and when the roles of the characters were unclear, the video was hard to follow. All these problems could be attributed to AI, and creators’ early over-reliance on the tool. Through conversations with the editor and each other, they were able to overcome this. It seemed to be correlated with both more group and editorial critique, as well as students having a better mental model of the strengths and weaknesses of the tool through repeated usage. The editor pointed out that “at some point [the students] were able to anticipate what the tool was going to suggest (to a point) because they understood how it was translating between genres [web stories to reels].”
@@
In the early videos, students often “miss the heart of the story.” For example, in a video about falling gas prices, the student “missed the main connection between the winter months and the change in gas prices”. Many exchanges in the early videos are wordy and contentless and do not help convey the news. It was important for students to understand the “news value” so the videos focused on the information rather than entertainment. The editor said that the students’ “learning/improvement happened when they made connections between the tool and their occupational standards.” Balancing between news and entertainment was one of the goals of the tool, and whereas it did help, it was still important for the creators to have their own understanding of this, so they could make good edits to what the tool was suggesting.
4.4.2. Learning from market research
In week 10 of the newsroom study, the team was given a report from the market research team detailing ways their news reels could be improved. This included more attention-grabbing elements early in the video, and more character motion-like actions within the scene (Section 5). These suggestions provided valuable insights because these were all emerging trends appearing in non-news-based social media videos that had not yet become integrated with news reels and were imbued into ReelFramer. Since news reels have to compete for attention with all the other reels, they will eventually have to adopt these trends to maintain their audience.
Working as a team, the creators brainstormed relevant actions they could put in their current videos and found several compelling ways to do this. For an article about tapping maple syrup trees early due to global warming, the video started with the character pouring syrup on waffles while talking. Another video about the opening of baseball season had several characters using props and doing simple, yet interesting actions like unwrapping food. A third added fun animation of a truck for a story about a monster truck rally.
The market research feedback was insightful, relevant and easy to apply. It was also outside the scope of AI, which cannot keep up with emerging trends.
5. Study 2 Findings: Audience Reception
Figure 4. Violin plots showing the distribution of ratings of videos as entertaining on a 1-7 scale. P is the professionally-created video, S the student-created video without using the tool and RF1-4, four videos created using the tool. The small letters labelling the bars identify groups of videos that are not statistically significantly different according to Tukey’s HSD. Multiple letters indicate that video is in two groups. Specifically, there is not a significant difference in the rating between videos P, S, RF2 and RF4, nor between video S and RF1-4 but the differences between video P and RF1 and RF3 are significant.
Figure 5. Violin plots showing the distribution of ratings of videos as entertaining. Labels are the same as Figure 4.
The main challenge in a news reel is to balance information and entertainment. We test how informative and entertaining news reels created in the study (using ReelFramer) were, compared to a popular professional news reel, and a news reel produced by a student at the same university but who was not in the study and was not exposed to ReelFramer.
First, we assessed how entertaining the news reels made during the story with ReelFramer were compared to the two baseline reels. According to a one-way ANOVA test, there was a significant difference in how entertaining the videos were ($F(5,229)=3.3$, $p=0.007$). Half of the videos produced using the tool were rated as as entertaining as the professionally made video. The non-study student’s video was also as entertaining as the popular video (Figure 4). However, it was rated less information than 3 of 4 ReelFramer videos.
Next, we assess how informative the news reels produced in the study are to the popular video and the non-study student reel. A one-way ANOVA test (not assuming equal variance in the groups) shows participant ratings of the video’s informativeness differed significantly ($F(5,108.18)=7.7$, $p<0.001$). Three-quarters of news reels produced in the study using ReelFramer were rated as “as informative as” the professionally produced video (Figure 5).
These results indicate that ReelFramer reels successfully balanced information and entertainment much of the time. They were competitive with the popular professional news reel, and it was better than the non-study student reel because they were more informational.
To assess audience acceptance of the news reels, we asked subjects whether they would watch each of the videos. Roughly three-quarters of subjects said they would watch the professionally produced video, 56% would watch the videos created with the tool and 45% would watch the student video produced without the tool. This difference was statistically significant ($\chi^{2}(5)=14.6$, $p=0.01$). Roughly half said they would share the professionally produced video, 40% would share the study videos, and only a quarter would the non-study student video. This difference was statistically significant ($\chi^{2}(5)=11.6$, $p=0.04$) There was no statistically significant difference in the perceived “fit” of videos on participant feeds ($\chi^{2}(5)=2.52$, $p=0.77$), though the student-produced video had a lower fit than other videos. Qualitative data suggests that this has to do with the subjects they prefer to watch (pets, sports, etc.) rather than quality. These results suggest that the videos produced with ReelFramer were closer to the professionally produced video in terms of acceptability than the non-study student video.
After conducting audience reception interviews, researchers identified strategic and manageable production changes they suggested to student journalists. Audience members indicated they would like more attention-grabbing elements early in the video and adding moving character movement within the scene. Student journalists wove this feedback into their subsequent productions and asked for more from the tool to match audience expectations. Similar to the way a newsroom would operate, journalists had to option of bringing market research into their production, continuing to iterate on their creative process.
6. Study 3 Findings: Professional Journalists
To supplement the results of the newsroom study done with student journalists, we conducted a 4-week study with two professional TV journalists from local news stations in the US.
6.1. Reasons for AI Tool Usage
Both journalists made 3 news reels each, for a total of 6 newsreels. They both chose to use the AI tool for all their videos. Similar to the newsroom study, they used the AI tool for framing and script writing, but not for storyboarding—they described framing and script writing as feeling “very different” than their training, and difficult to “get started with.” They also said that they had to edit “about half” the AI scripts for clarity, correctness, and smooth language.
6.2. Role of Professional Human Creativity: Human Creativity Where AI Leaves off
In four of the six videos, the journalists made notable creative additions. This included some of the same categories seen in the student study—adding personality to characters through acting and adding visual effects. However, they also introduced new elements. Two videos included real footage as background imagery. One video used an audio track. One video took the role-play premise to an extreme and introduced a time-lapse—where a citizen called city services over the course of 10 years to get a fallen tree removed (the city representative promised to “get back to them” in a robotic voice in every call). With professional journalism training and resources, it is possible to add even more creativity where the AI tool leaves off.
Of the six videos in the study, none experienced the major AI failures that forced students to make the news without the tool. However, AI did create some problems. Participant Pro 1 made a video discussing a new scam. The character introducing the scam was the Better Business Bureau (BBB). Although the typical TV news audiences know what the BBB is, Gen-Z and even Millennial audiences are less familiar with and are confused about the connection between the BBB and scams. AI was not aware of this gap in audience knowledge, but the editor was. But that decision still required much discussion and consideration.
6.3. AI Risks: Even Professionals Need to Hone Critique of AI
Major Issues (6) AI Attributed (4) Non-AI Attributed(2) unnecessary dialog (2) slow paced (1) too few facts (1) added timeline narrative unclear (1) too long (1)
Table 4. Major Issues: AI vs Non-AI Attributed issues in news reels created in the professional study.
Minor Issues (8) AI Attributed (5) Non-AI Attributed(3) confusing character names (3) distracting music (2) awkward AI phrase (2) should not name specific unrelated brands (1)
Table 5. Minor Issues: AI vs Non-AI Attributed issues in news reels created in the professional study.
Despite years of journalistic practice, the professionals still made many errors. Four of the six major errors and five of eight minor errors were introduced by AI and not caught by professionals before filming (see Tables 4 and 5). Generally, they were the same types of errors seen in the student study. This included major errors like having unnecessary and contentless dialog (“I can’t thank you enough”, “We’re so happy to help”), having too few facts, and being too long. As well as minor issues, like using awkward AI phrases (“How about I come and start training my own railings?”). The editor thought these should have been glaring errors to catch. Upon viewing the iteration of the video containing that dialog, the editor asked if “training railings” was a concept the professional journalist knew (perhaps a local phrase). The journalist said no, “The AI said it; I thought it was a thing.” Despite years of journalism training—a field that prides itself on being critical—these professionals were imperfect AI critics.
6.4. Professional Creative Role Over Time
Similar to the student study, news reels made by professional journalists still required major edits. However, both professionals required no edits by their third video (see Table 6). In both videos needing major edits, the errors were similar to the students’—the news content was unclear due to a variety of errors, like contentless dialog, and unclear roles. They did have to be coached by the editor to correct AI, but once the editor pointed that out, they adapted almost immediately. This is likely because they had a wealth of journalism training and understanding of “news value”, and thus faster to integrate their news training into their use of the tool.
Participant Article 1 Article 2 Article 3 Pro1 Major Minor No edits AI errors 1 3 Non-AI errors 3 1 Pro2 Minor Major No edits AI errors 1 3 Non-AI errors 1
Table 6. Classification of the amount of editing needed for news reels made by professional journalists.
7. Discussion and Conclusion
7.1. Experts Value AI Creative Support and Use Their Creativity Where AI Leaves Off
We found that when AI presents a solution, that solution, unless obviously flawed, is generally adopted. The user then applies their creativity to the parts of the project not solved by AI. There may be something to be learned from other areas of design. Lars Lerup, in a book called Building the Unfinished, observed that the physical environment will change over time: for example, we renovate buildings (Lerup, 1977). Instead of designing something that is a finished product, it might be better to build it in a way that encourages or facilitates later modification. A contemporaneous and related idea from architecture has to do with planning and building incrementally rather than all at once (Alexander, 1977).
AI leaves many things unfinished. In the newsroom study, AI failed in spectacular ways, leaving people to complete the task. AI also does not know what the next trend will be. In the newsroom study, we saw how the results from the market research were immediately used by journalists to keep up with social media engagement trends that AI was unaware of. Creativity is about discovering what comes next, and so AI has not shown abilities to do that.
With all automated tools there is a fear that human skills will atrophy from disuse. However, in this study, we saw rampant use of creativity, especially where the tool left off. This incompleteness of AI could be thought of as a feature rather than a bug—it is an opportunity to keep us engaged with the tool. Future work might explore the different ways of leaving an artifact unfinished and their effects on the quality of the end product, the overall level of productivity, and the long-term effect on human skill.
7.2. Creating Vigilance to AI Errors
AI makes errors. Often it makes obvious errors like using totally made up phrases. It was surprising to see that journalists, even professionals, filmed videos without fixing them. There seems to be a pervasive assumption that AI is an expert, even when AI is operating in a field in which the human is the expert and the AI is meant to play a supporting role.
Whatever the reason for their deference, people need to develop a vigilance towards AI errors. AI is so easy to generate that it seems to lull us into System I reasoning (Kahneman, 2011), where we think fast and mostly go with the flow. However, when working with AI, we need to develop ways to help people shift into System II reasoning, where we think slowly and think critically. System I is our natural state because it will take more than just willpower to develop this vigilance.
Teams and organizations will have to design ways to avoid falling for AI errors. Editors are needed to be critical; editors will need to learn the new flaws that AI is introducing into creative work. For example, AI is known to lean toward cliches in creative domains  (Chakrabarty et al., 2024; Mirowski et al., 2024), known to make up phrases (as shown here), known to lack social context, and known to be bad at simple math (Schick et al., 2023). When possible, it might also be helpful to constantly track what text or images were AI-generated so editors know where to focus their attention.
@@ 7.3. Limitations and Future Work
These studies attempt to explore AI at work in a newsroom, but we only studied one news team—and it was one that we assembled for this purpose. Although it had many of the traits of a real small-town newsroom, there are limitations in how much we can generalize to all newsrooms. Although the study lasted 14 weeks, there is no telling what would have happened if the study had continued after those 14 weeks. While we did not see any skill atrophy in the 14-week period, perhaps we would have after a year.
This paper focused on one particular tool for one particular social media phenomenon, but different tools and trends may have different benefits and challenges. However, factors such as accepting AI errors seem to be pervasive problems. ReelFramer as a tool is certainly not perfect. If its workflow had been better, perhaps creators would not have had so many opportunities to be creative when it failed. However, the most important failures were the ones when the new reel form did not fit an article—which was not a property of the tool, but of the narrative form.
It would be interesting future work to see if people were still using the tool in a year or whether they had embodied the news reel form so well the tool no longer brought them cognitive benefit. Or perhaps they might even find it too limiting over time, as they started to understand the design space and perhaps start to innovate on it. It would also be interesting to study how the team might improve the tool for themselves. Since the tool is basically a structured prompting technique, the creators could easily edit the tool by simply editing the prompts in the text. Another form of creativity would be to use prompts within the tool to innovate on the form.

### v2

@@ 3.4.1. Editors’ Reflection Over Time
Through conversations with the editor and each other, creators were able to overcome this. It seemed to be correlated with both more group and editorial critique, as well as students having a better mental model of the strengths and weaknesses of the tool through repeated usage. The editor pointed out that “at some point [the students] were able to anticipate what the tool was going to suggest (to a point) because they understood how it was translating between genres [web stories to reels].”
@@
In the early videos, students often “miss the heart of the story.” It was important for students to understand the “news value,” so the videos focused on the information rather than entertainment. The editor said that the students’ “learning and improvement happened when they made connections between the tool and their occupational standards.” Balancing between news and entertainment was one of the goals of the tool, and whereas it did help, it was still important for the creators to have their own understanding of this, so they could make appropriate edits.
@@ 3.5. Limitations and Future Work
This study examined one student newsroom using a specific AI tool for 14 weeks, limiting generalizability. Future work could explore longer-term effects on populations of multiple types of journalistic and AI-based experiences. In particular, professional newsrooms may face different dynamics given their tighter deadlines, higher editorial standards, and larger audiences, and other creative fields may reveal distinct patterns of human-AI collaboration.
4. Conclusion
This paper studied the benefits and challenges of using AI in the organizational context of a newsroom. AI has many strengths as a creativity support tool—especially for writing—so we studied journalists, experts in the craft of writing and storytelling. Newsrooms and other creative industries are places where AI could significantly impact work—either by changing its nature or potentially displacing it. Overall, we found that AI was helpful both as scaffolding during the learning phase and as a creative springboard once users understood the task. AI supported human creators but did not replace them. While it could generate outputs, it was the creators who took ownership of the task, using judgment and creativity to tailor stories for audiences, rather than merely adapting them to templates. Despite the flexibility of generative AI, it follows instructions linearly and struggles to adapt to the nuances of individual stories or audiences. Because AI still makes frequent mistakes and lacks contextual understanding, editorial oversight was essential to ensure quality and avoid blindly accepting flawed outputs. Through this reflective process, creators also reasserted their expertise. While AI was initially seen as an “authority,” users ultimately learned to balance its contributions with their own.